# Neural Topic Modeling with large language models in the loop

### v1

@@ Appendix F Confidence Alternatives
Figure F1: Study of confidence alternatives. Learning curves of LLM-ITL (ETM) with different confidence in terms of topic coherence (NPMI) and topic alignment (PN) on 20News ((a) and (b)) and DBpedia ((c) and (d)). Error bars are omitted for clarity in the figure.
@@
Considering the efficiency of including confidence scores during the training of LLM-ITL, we focus solely on single-sample approaches for LLM uncertainty estimation in our study, where we run a single round of LLM inference for a given topic. We consider the following confidence alternatives to our topic labeling confidence for our study: (1) No Conf: No LLM confidence estimation is included, and $\text{Conf}(\bm{w}_{k}^{l})=1$ in Eq. 10 during the training. (2) Verbalized confidence (Xiong et al., 2023), which directly asks the LLM for its confidence in solving a problem. The prompt we used for eliciting verbalized confidence is shown in Figure A2. (3) Self-Reflective confidence (Chen & Mueller, 2023), which prompts the LLM to evaluate its own answer in a two-stage manner. The topic label is obtained in the first stage, and the LLM evaluates this answer in a follow-up question (Figure A3). (4) p(True) (Kadavath et al., 2022), which is similar to self-reflective confidence, but asks a true/false question instead. It takes the token probability of the response as the confidence measure (Figure A3). (5) SeqLike (Ren et al., 2022), which computes the length-normalized sequence likelihood of the output from the LLM.
@@
From the results in Figure F1, we can observe that while verbalized confidence improves topic coherence better, it biases the topics towards the LLM’s knowledge rather than the topics within the corpus, leading to reduced topic alignment. On the other hand, label token probability and word intrusion confidence consistently yield the best topic alignment performance, suggesting greater relevance of the topics to the corpus.
Appendix G Hyper-parameter Studies
@@
Figure G1: Learning curves of LLM-ITL (ETM) with different $\gamma$ in terms of NPMI and PN on 20News ((a) and (b)) and DBpedia ((c) and (d)).
Here, we study the hyper-parameter of LLM-ITL, focusing on topic refinement strength $\gamma$. We vary its value from 1 to 400 and plot the learning curves in terms of NPMI and PN, as shown in Figure G1. We can observe that: (1) In terms of topic coherence, $\gamma$ values between 100 and 300 yield similar performance, suggesting low sensitivity to $\gamma$ within this range. (2) In terms of topic alignment, a higher $\gamma$ leads to slightly reduced performance in the initial phase. This occurs because relying heavily on topic refinement from the LLM causes the topics to bias towards the LLM’s knowledge rather than the information from the input corpus. However, as training progresses, the performance converges to similar values. These observations suggest low sensitivity to $\gamma$ and its flexibility in controlling the balance between the information from the corpus and the knowledge from the LLM.
@@
Figure G2: Learning curves of LLM-ITL (ETM) with different settings for the number of words (i.e., $N$) in the topic label, evaluated in terms of NPMI and PN on 20News ((a) and (b)) and DBpedia ((c) and (d)).
@@
Here, we examine the hyper-parameter within the prompt, i.e., the number of words $N$ used for the topic label. We vary the number of words for the topic label from 1 to 5 and plot the learning curves in terms of NPMI and PN. As illustrated in Figure G2, we can observe that: (1) Using more words, such as a 5-word topic label (e.g., $N=5$), results in the least improvement in topic coherence, while using a 2-word topic label (e.g., $N=2$) achieves the best performance. (2) As for topic alignment performance, the number of words in the topic label shows comparable performance.
@@ Appendix H Study of prompts
Table H1: Prompt variants performance. The best and second-best performance of each column are highlighted in boldface and underlined, respectively.
Origin 0.978 224.48 197.86 7.165
Variant_1 0.967 228.48 80.81 4.520
Variant_2 0.935 269.48 167.05 3.720
Variant_3 0.980 223.48 172.51 6.146
Variant_4 0.972 189.48 191.04 7.490
Variant_5 0.956 208.48 159.66 6.215
Iterative Refinement (Chang et al., 2024) 0.993 1872.13 936.70 2.845
Figure H1: Learning curves of LLM-ITL (ETM) with different prompts in terms of NPMI and PN on 20News (figure (a) and (b)) and DBpedia (figure (c) and (d)).
@@
Here, we study the effectiveness of different prompts for topic suggestion. We obtain variants of topic suggestion prompts in Figure A1 through modifying the technique used in PromptBreeder (Fernando et al., 2023). To be specific, we first create a set of 100 ‘mutation-prompts’ (e.g., “Make a variant of the prompt”) and 100 ‘thinking-styles’ (e.g., “Let’s think step by step”). We generate a set of 50 task prompts by concatenating a randomly drawn ‘mutation-prompt’ and a randomly drawn ‘thinking-style’ to the original prompt, and provide that to the Claude 3.5191919https://www.anthropic.com/news/claude-3-family to produce a continuation, resulting in a different task prompt. Secondly, we randomly select 50 topics over 4 experiment datasets. We run those topics through 50 generated task prompts and filter out the generated prompts that cannot give JSON format in the selected topics or generate above 300 tokens. We are left with 14 topics. We then leverage Claude 3.5 to judge the quality of generated topics and refined topic words. We rank 14 methods by overall topics and refined topic words to get 5 variants of prompts. In addition to the prompt variants generated by those steps, we also investigate the topic refinement prompt used in Chang et al. (2024) (see Figure 2 of their paper). All the prompt variants for topic refinement in this study are illustrated in Table H2.
@@
We randomly sample 1000 topics learned by topic models, then use different prompts to refine the topics with LLAMA3-8B-Instruct [^20] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct. We analyze the effectiveness of prompts in different aspects, including Success Rate (Ulmer et al., 2024): the proportion of cases where the target answer can be successfully extracted from the LLM’s output; N_Input and N_Output (Chang et al., 2024): the average number of tokens of input and output of the LLM; and Refined TC: the average NPMI scores of the refined topics.
@@ Results
Based on the above observations, we further investigate the effectiveness of the improved prompt within the LLM-ITL framework. We plot the learning curves of LLM-ITL using the original prompt and its variant (Variant_4, which shows better performance from Table H1). We observe that the overall performance in terms of both metrics is comparable.
@@
Prompt
Origin Analyze step-by-step and provide the final answer. Step 1. Given a set of words, summarize a topic (avoid using proper nouns as topics) by 2 words that covers most of those words. Note, only the topic, no other explanations. Step 2. Remove irrelevant words about the topic from the given word list. Note, only the removed words, no other explanations. Step 3. Add new relevant words (maximum 10 words) about the topic to the word list up to 10 words. Note, only the added words, no other explanations. Step 4. Provide your answer in json format as {‘Topic’: ‘<2 Word Topic>’, ‘Words’: ‘<Refined 10 Word List>’}. Note, only 10 refined words allowed for the topic, and no follow up explanations.
Variant_1 Perform the following actions sequentially and provide the final result: Step 1. After examining a set of words, condense a subject (avoid proper nouns) into 2 words that encompass most of those words. (Note: Only the subject, no further elaboration.) Step 2. Eliminate irrelevant words from the given word list based on the subject. (Note: Only the removed words, no further elaboration.) Step 3. Add new pertinent words (maximum 10 words) related to the subject to the word list until it reaches 10 words. (Note: Only the added words, no further elaboration.) Step 4. Present your response in JSON format as {‘Topic’: ‘<2 Word Subject>’, ‘Words’: ‘<Refined 10 Word List>’}. Note: Only 10 refined words are permitted for the subject, and no follow-up explanations.
Variant_2 Perform a meticulous examination and furnish the conclusive resolution. Stride 1. Bestowed a catalogue of vocabularies, condense a subject matter (circumvent the employment of proper appellations as subjects) by dual words that envelop the preponderance of those vocabularies. (Heed, solely the subject, devoid of supplemental explication.) Stride 2. Dislodge irrelevant vocabularies concerning the subject from the granted vocabulary catalogue. (Heed, solely the dislodged vocabularies, devoid of supplemental explication.) Stride 3. Amalgamate novel applicable vocabularies (maximal 10 vocabularies) concerning the subject to the vocabulary catalogue up to 10 vocabularies. (Heed, solely the amalgamated vocabularies, devoid of supplemental explication.) Stride 4. Tender your resolution in json format as {‘Topic’: ‘<2 Word Subject>’, ‘Words’: ‘<Refined 10 Word Catalogue>’}. Heed, solely 10 refined vocabularies permitted for the subject, and devoid of successive explication.
Variant_3 Step-by-step analysis and final answer: Step 1. Given a set of words, summarize a topic (avoid using proper nouns as topics) by 2 words that covers most of those words. (Note, only the topic, no other explanations.) Step 2. Remove irrelevant words about the topic from the given word list. (Note, only the removed words, no other explanations.) Step 3. Add new relevant words (maximum 10 words) about the topic to the word list, keeping the total word count at 10 words. (Note, only the added words, no other explanations.) Step 4. Provide your answer in JSON format as {‘Topic’: ‘<2 Word Topic>’, ‘Words’: ‘<Refined 10 Word List>’}. Note, only 10 refined words allowed for the topic, and no follow-up explanations.
Variant_4 Break down the analysis into steps and give the final response. 1. Look at a set of words and identify a 2-word topic that sums up most of those words (don’t use proper nouns as topics, just state the topic). 2. Remove words from the list that don’t relate to the topic (just list the removed words). 3. Add new relevant words about the topic to the list, up to 10 words total (just list the new added words). 4. Provide your response in JSON format: {‘Topic’: ‘<2 Word Topic>’, ‘Words’: ‘<Refined 10 Word List>’}. Only include 10 words for the refined list, no explanations.
Variant_5 Step-by-step analysis and provide the final answer in JSON format: Step 1: Based on the given set of words, summarize a topic using 2 words that encompass most of those words (avoid proper nouns). Step 2: Remove any irrelevant words from the given word list that do not relate to the summarized topic. Step 3: Add new relevant words (up to 10 words) that are related to the summarized topic. Step 4: Present your answer in the following JSON format: {‘Topic’: ‘<2 Word Topic>’, ‘Words’: ‘<Refined 10 Word List>’}, where ‘Topic’ contains the 2-word summarized topic, and ‘Words’ contains the refined list of 10 words related to that topic. Do not provide any additional explanations.

### v2

@@ Appendix F Study of Confidence Alternatives
Figure F1: Learning curves of LLM-ITL (ETM as the base model) with different LLM confidences in terms of topic coherence ($C_{V}$) and topic alignment (PN) on 20News. Error bars are omitted for clarity in the figure.
@@
Here, we compare our proposed topic labeling confidence with other LLM confidence alternatives within the LLM-ITL framework. We limit our focus to single-sample approaches, where a single round of LLM inference is performed for a given topic. This ensures they can be efficiently applied within the training of LLM-ITL. We consider the following confidence alternatives to our topic labeling confidence in this study: (1) No Conf., where no LLM confidence estimation is included, and $\text{Conf}(\bm{w}_{k}^{l})=1$ in Eq. 10 during the training. (2) Verbalized Confidence (Xiong et al., 2023), which directly asks the LLM for its confidence in solving a problem. The prompt we used for eliciting verbalized confidence is shown in Figure A2. (3) Self-Reflective Confidence (Chen and Mueller, 2023), which prompts the LLM to evaluate its own answer in a two-stage manner. The topic label is obtained in the first stage, and the LLM evaluates this answer in a follow-up question (Figure A3). (4) p(True) (Kadavath et al., 2022), which is similar to self-reflective confidence, but asks a true/false question instead. It takes the token probability of the response as the confidence measure (Figure A3). (5) SeqLike. (Ren et al., 2022), which computes the length-normalized sequence likelihood of the output from the LLM.
@@
From the results in Figure F1, we can observe that while verbalized confidence results in greater improvements in topic coherence, it can bias the topics toward the LLM’s knowledge rather than the input corpus, leading to reduced topic alignment. In contrast, label token probability and word intrusion confidence consistently achieve the best topic alignment performance, indicating a stronger relevance of the topics to the corpus.
Appendix G Study of Hyper-parameters
@@
Figure G1: Learning curves of LLM-ITL (ETM as the base model) with different $\gamma$ in terms of topic coherence ($C_{V}$) and topic alignment (PN) on 20News.
@@
Figure G2: Learning curves of LLM-ITL (ETM as the base model) with different settings of $N$ (i.e., the number of words in the topic label) in terms of topic coherence ($C_{V}$) and topic alignment (PN) on 20News.
Here, we study the hyper-parameter of LLM-ITL, focusing on topic refinement strength $\gamma$, and the number of words $N$ used for the topic label.
As for $\gamma$, we vary its value from 1 to 400 and plot the learning curves in terms of $C_{V}$ and PN, as shown in Figure G1. We can observe that: (1) In terms of topic coherence, $\gamma$ values between 100 and 300 yield similar performance, suggesting low sensitivity to $\gamma$ within this range. (2) In terms of topic alignment, a higher $\gamma$ leads to slightly reduced performance in the initial phase. This occurs because relying heavily on topic refinement from the LLM causes the topics to bias towards the LLM’s knowledge rather than the information from the input corpus. However, as training progresses, they converge to similar values. These observations suggest that LLM-ITL exhibits low sensitivity to $\gamma$ within a certain range and offers flexibility in controlling the balance between learning from the corpus and the LLM.
Origin 0.978 224.48 197.86 0.554
Variant_1 0.967 228.48 80.81 0.512
Variant_2 0.935 269.48 167.05 0.498
Variant_3 0.980 223.48 172.51 0.537
Variant_4 0.972 189.48 191.04 0.562
Variant_5 0.956 208.48 159.66 0.543
Iterative Refinement (Chang et al., 2024) 0.993 1872.13 936.70 0.478
Table H1: Study of prompt variants. The best and second-best performance of each column are highlighted in boldface and underlined, respectively.
@@
As for $N$, we vary the number of words for the topic label from 1 to 5 and plot the learning curves in terms of $C_{V}$ and PN. As illustrated in Figure G2, we can observe that: (1) Using more words as the topic label for LLM-ITL, such as a 5-word topic label (e.g., $N=5$), results in the least improvement in topic coherence. While using a 2-word topic label (e.g., $N=2$) achieves the best performance in terms of topic coherence. (2) For topic alignment, the number of words in the topic label demonstrates comparable performance, suggesting low sensitivity to $N$.
@@ Appendix H Study of Prompts
Here, we study the effectiveness of different prompts for topic suggestion. We obtain variants of topic suggestion prompts in Figure A1 through modifying the technique used in PromptBreeder (Fernando et al., 2023). To be specific, we first create a set of 100 “mutation-prompts” (e.g., “Make a variant of the prompt") and 100 “thinking-styles” (e.g., “Let’s think step by step"). We generate a set of 50 task prompts by concatenating a randomly drawn “mutation-prompt” and a randomly drawn “thinking-style” to the original prompt, and provide that to the Claude 3.5101010https://www.anthropic.com/news/claude-3-family to produce a continuation, resulting in a different task prompt. Secondly, we randomly select 50 topics over 4 experiment datasets. We run those topics through 50 generated task prompts and filter out the generated prompts that cannot give JSON format in the selected topics or generate above 300 tokens. We are left with 14 topics. We then leverage Claude 3.5 to judge the quality of generated topics and refined topic words. We rank 14 methods by overall topics and refined topic words to get 5 variants of prompts. In addition to the prompt variants generated by those steps, we also investigate the topic refinement prompt used in Chang et al. (2024) (see Figure 2 of their paper). All the prompt variants for topic refinement in this study are illustrated in Table H2.
Figure H1: Learning curves of LLM-ITL (ETM as the base model) with different prompts in terms of $C_{V}$ and PN on 20News.
@@
We randomly sample 1000 topics learned by topic models, then use different prompts to refine the topics with LLAMA3-8B-Instruct. We analyze the effectiveness of prompts in different aspects, including Success Rate (Ulmer et al., 2024): the proportion of cases where the target answer can be successfully extracted from the LLM’s output; N_Input and N_Output (Chang et al., 2024): the average number of tokens of input and output of the LLM; and Refined TC: the average topic coherence scores of the refined topics.
@@ Results
Based on the above observations, we further investigate the effectiveness of the improved prompt within the LLM-ITL framework. We plot the learning curves of LLM-ITL using the original prompt and its variant (Variant_4, which shows better performance from Table H1). We observe that the overall performance in terms of both metrics is comparable by using both prompts.
Prompt
Origin Analyze step-by-step and provide the final answer. Step 1. Given a set of words, summarize a topic (avoid using proper nouns as topics) by 2 words that covers most of those words. Note, only the topic, no other explanations. Step 2. Remove irrelevant words about the topic from the given word list. Note, only the removed words, no other explanations. Step 3. Add new relevant words (maximum 10 words) about the topic to the word list up to 10 words. Note, only the added words, no other explanations. Step 4. Provide your answer in json format as {“Topic”: “<2 Word Topic>”, “Words”: “<Refined 10 Word List>”}. Note, only 10 refined words allowed for the topic, and no follow up explanations.
Variant_1 Perform the following actions sequentially and provide the final result: Step 1. After examining a set of words, condense a subject (avoid proper nouns) into 2 words that encompass most of those words. (Note: Only the subject, no further elaboration.) Step 2. Eliminate irrelevant words from the given word list based on the subject. (Note: Only the removed words, no further elaboration.) Step 3. Add new pertinent words (maximum 10 words) related to the subject to the word list until it reaches 10 words. (Note: Only the added words, no further elaboration.) Step 4. Present your response in JSON format as {“Topic”: “<2 Word Subject>”, “Words”: “<Refined 10 Word List>”}. Note: Only 10 refined words are permitted for the subject, and no follow-up explanations.
Variant_2 Perform a meticulous examination and furnish the conclusive resolution. Stride 1. Bestowed a catalogue of vocabularies, condense a subject matter (circumvent the employment of proper appellations as subjects) by dual words that envelop the preponderance of those vocabularies. (Heed, solely the subject, devoid of supplemental explication.) Stride 2. Dislodge irrelevant vocabularies concerning the subject from the granted vocabulary catalogue. (Heed, solely the dislodged vocabularies, devoid of supplemental explication.) Stride 3. Amalgamate novel applicable vocabularies (maximal 10 vocabularies) concerning the subject to the vocabulary catalogue up to 10 vocabularies. (Heed, solely the amalgamated vocabularies, devoid of supplemental explication.) Stride 4. Tender your resolution in json format as {“Topic”: “<2 Word Subject>”, “Words”: “<Refined 10 Word Catalogue>”}. Heed, solely 10 refined vocabularies permitted for the subject, and devoid of successive explication.
Variant_3 Step-by-step analysis and final answer: Step 1. Given a set of words, summarize a topic (avoid using proper nouns as topics) by 2 words that covers most of those words. (Note, only the topic, no other explanations.) Step 2. Remove irrelevant words about the topic from the given word list. (Note, only the removed words, no other explanations.) Step 3. Add new relevant words (maximum 10 words) about the topic to the word list, keeping the total word count at 10 words. (Note, only the added words, no other explanations.) Step 4. Provide your answer in JSON format as {“Topic”: “<2 Word Topic>”, “Words”: “<Refined 10 Word List>”}. Note, only 10 refined words allowed for the topic, and no follow-up explanations.
Variant_4 Break down the analysis into steps and give the final response. 1. Look at a set of words and identify a 2-word topic that sums up most of those words (don’t use proper nouns as topics, just state the topic). 2. Remove words from the list that don’t relate to the topic (just list the removed words). 3. Add new relevant words about the topic to the list, up to 10 words total (just list the new added words). 4. Provide your response in JSON format: {“Topic”: “<2 Word Topic>”, “Words”: “<Refined 10 Word List>”}. Only include 10 words for the refined list, no explanations.
Variant_5 Step-by-step analysis and provide the final answer in JSON format: Step 1: Based on the given set of words, summarize a topic using 2 words that encompass most of those words (avoid proper nouns). Step 2: Remove any irrelevant words from the given word list that do not relate to the summarized topic. Step 3: Add new relevant words (up to 10 words) that are related to the summarized topic. Step 4: Present your answer in the following JSON format: {“Topic”: “<2 Word Topic>”, “Words”: “<Refined 10 Word List>”}, where “Topic” contains the 2-word summarized topic, and “Words” contains the refined list of 10 words related to that topic. Do not provide any additional explanations.